a short run production function is given as, Q=L3- 2L2 - 4L+40 determine the number of laborers that should be used to maximize total product



1
Expert's answer
2022-02-10T09:45:33-0500

To maximize total product we have to solve the equation:

"Q'=(L^3-2L^2-4L+40)'=0"

"3L^2-4L-4=0" (SOLVE THE QUADRATIC EQUATION)

"L_1=2"

"L_2=-0.6667"